# A General Class of Model-Free Dense Precision Matrix Estimators

### v1

@@ 6 Conclusion
In this appendix, we only cover the proof of Theorem 2, the rest of the proofs are relegated to Online Appendix. Before the following proof, we proceed with three inequalities that are used for rectangular matrices. Take a generic rectangular matrix $M$ and a vector $x$.
@@
where .$\|.\|_{F}$ is the Frobenius norm and we get the first inequality by (4.4) of Vershynin (2019).
@@
(i). Note that we need to estimate the $j$ th main diagonal term for the precision matrix of asset returns. In that sense the key element is the reciprocal of that term which is defined in (12), and the population version in (OB.11). We use triangle inequality to consider the estimation error by (12)(OB.11)
@@
where $C_{\psi}$ are the upper bound constant for the subexponential norm in Definition 2.7.5 of Vershynin (2019). Next $2p^{2}exp\left(-cmin(\frac{t_{1}}{C_{\psi}^{2}},\frac{t_{1}}{C_{\psi}})n\right)=exp\left(log(2p^{2})-c\frac{t_{1}^{2}n}{C_{\psi}^{2}}\right)$, with $t_{1}:=C\frac{\sqrt{logp}}{\sqrt{n}}$ and with sufficiently large $n$ (i.e.$\sqrt{n}\geq\frac{C}{C_{\psi}}(logp)^{1/2}$). Set $c_{\psi}:=c\frac{C^{2}}{C_{\psi}^{2}}-2>0$ to have $exp[log2+2logp-c\frac{C^{2}}{C_{\psi}^{2}}logp]=\frac{2}{p^{c_{\psi}}}.$ Then
@@
The asymptotics in (A.14) are obtained after taking $n\to\infty,p\to\infty$. Consider the third right side term in (A.9). Lemma 2.7.7 and Corollary 2.8.3 of Vershynin (2019) together with Assumptions 1, 4, and Holder’s inequality give
@@
where the last equality uses $p-1>K$. By (A.13)-(A.15) we have the first right side term in (A.8)
@@
Use these last two equations
$\max_{1\leq j\leq p}\|\frac{1}{n}\sum_{t=1}^{n}[y_{jt}Y_{-jt}^{\prime}-E(y_{jt}-\mu_{j})(Y_{-jt}-\mu_{-j})^{\prime}]\|_{2}\leq\max_{1\leq j\leq p}\|a_{j}^{\prime}\left(\frac{1}{n}\sum_{t=1}^{n}[f_{t}f_{t}^{\prime}-\Sigma_{f}]\right)A_{-j}^{\prime}\|_{2}$ (A.17)
$+$ $\max_{1\leq j\leq p}\|(\frac{1}{n}\sum_{t=1}^{n}u_{jt}f_{t}^{\prime})A_{-j}^{\prime}\|_{2}+\max_{1\leq j\leq p}\|a_{j}^{\prime}(\frac{1}{n}\sum_{t=1}^{n}f_{t}U_{-jt}^{\prime})\|_{2}+\max_{1\leq j\leq p}\|\frac{1}{n}\sum_{t=1}^{n}u_{jt}U_{-jt}^{\prime}\|_{2}.$
@@
$=$ $O(K^{1/2})O_{p}(\max(\frac{K^{1/2}}{n^{1/2}},\frac{\sqrt{logn}}{n^{1/2}}))O(d_{1n})=O_{p}(max(\frac{Kd_{1n}}{n^{1/2}},\frac{d_{1n}\sqrt{Klogn}}{n^{1/2}})),$
@@
$=$ $O_{p}(\sqrt{K}\frac{\sqrt{logp}}{\sqrt{n}})O(d_{1n}),$
where we use (A.1) for the first inequality since $u_{jt}f_{t}^{\prime}:1\times K$ vector and $A_{-j}^{\prime}:K\times p-1$ matrix, and the second inequality is by simple vector norm inequality tying $l_{2},l_{\infty}$ vector norms, and the rates are by Assumption 3(iii) $\max_{1\leq j\leq p}\|A_{-j}^{\prime}\|_{2}=O(d_{1n})$ and Lemma 2.7.7 and Corollary 2.8.3. of Vershynin (2019) which provides $\|n^{-1}\sum_{t=1}^{n}u_{jt}f_{t}^{\prime}\|_{\infty}=O_{p}(\sqrt{logp/n})$. Now third term on the right side of (A.17)
@@
$r_{w1}:=\max\left(\sqrt{\frac{logn+\bar{r}}{n\bar{\eta}}},\sqrt{\frac{K}{\bar{\xi}\delta_{n}}},\sqrt{\frac{K\bar{\Psi}}{\bar{\xi}\bar{\eta}\delta_{n}}},\sqrt{\frac{K}{\bar{\xi}\bar{\eta}}}\right),$ (A.23)
The proof of Theorem 1-Step 2d together with $r_{n}/\delta_{n}\to 1$, gives
@@
By (A.26), the third term on the right side of (A.8) is by Theorem 1
$\max_{1\leq j\leq p}\|E(y_{jt}-\mu_{j})(Y_{-jt}-\mu_{-j})^{\prime}\|_{2}\max_{1\leq j\leq p}\|\hat{\alpha}_{j}-\alpha_{j}^{*}\|_{2}=O_{p}(d_{2n}r_{w2}),$ (A.27)
with $r_{w2}:=\max(r_{w1},\sqrt{\frac{K}{\bar{\xi}}})$.
@@
$\max_{1\leq j\leq p}E[(y_{jt}-\mu_{j})(Y_{-jt}-\mu_{-j})^{\prime}]\alpha_{j}^{*}$$\leq$ $\left[\max_{1\leq j\leq p}\|E[(y_{jt}-\mu_{j})(Y_{-jt}-\mu_{-j})^{\prime}]\|_{2}\right]\left[\max_{1\leq j\leq p}\|\alpha_{j}^{*}\|_{2}\right]$ (A.28)

### v2

@@ 6 Conclusion
Before the following proof, we proceed with three inequalities that are used for rectangular matrices. Take a generic rectangular matrix $M$ and a vector $x$.
@@
where $\|.\|_{F}$ is the Frobenius norm and we get the first inequality by (4.4) of Vershynin (2019).
@@
(i). Note that we need to estimate the $j$ th main diagonal term for the precision matrix of asset returns. In that sense the key element is the reciprocal of that term which is defined in (11), and the population version in (B.10). We use triangle inequality to consider the estimation error by (11)(B.10)
@@
[…]
@@
The asymptotics in (A.101) are obtained after taking $n\to\infty,p\to\infty$. Consider the third right side term in (A.96). Lemma 2.7.7 and Corollary 2.8.3 of Vershynin (2019) together with Assumptions 1, 4 and Holder’s inequality give
@@
where the last equality uses p1>Kp-1>K. By (A.100)-(A.102) we have the first right side term in (A.95)
@@
Use these last two equations to see that the following term
$\max_{1\leq j\leq p}\|\frac{1}{n}\sum_{t=1}^{n}[y_{jt}Y_{-jt}^{\prime}-E(y_{jt}-\mu_{j})(Y_{-jt}-\mu_{-j})^{\prime}]\|_{2},$
is upper bounded by
$\max_{1\leq j\leq p}\|a_{j}^{\prime}\left(\frac{1}{n}\sum_{t=1}^{n}[f_{t}f_{t}^{\prime}-\Sigma_{f}]\right)A_{-j}^{\prime}\|_{2}+\max_{1\leq j\leq p}\|(\frac{1}{n}\sum_{t=1}^{n}u_{jt}f_{t}^{\prime})A_{-j}^{\prime}\|_{2}$ (A.104)
$+$ $\max_{1\leq j\leq p}\|a_{j}^{\prime}(\frac{1}{n}\sum_{t=1}^{n}f_{t}U_{-jt}^{\prime})\|_{2}+\max_{1\leq j\leq p}\|\frac{1}{n}\sum_{t=1}^{n}u_{jt}U_{-jt}^{\prime}\|_{2}.$
@@
$=$ $O(K^{1/2})O_{p}(\max(\frac{K^{1/2}}{n^{1/2}},\frac{\sqrt{logn}}{n^{1/2}}))O(\sqrt{p})=O_{p}(max(\frac{K\sqrt{p}}{n^{1/2}},\frac{\sqrt{p}\sqrt{Klogn}}{n^{1/2}})),$
@@
$=$ $O_{p}(\sqrt{K}\frac{\sqrt{logp}}{\sqrt{n}})O(\sqrt{p}),$
where we use (A.88) for the first inequality since $u_{jt}f_{t}^{\prime}:1\times K$ vector and $A_{-j}^{\prime}:K\times p-1$ matrix, and the second inequality is by simple vector norm inequality tying $l_{2},l_{\infty}$ vector norms, and the rates are by Assumption 3(iii) $\max_{1\leq j\leq p}\|A_{-j}^{\prime}\|_{2}=O(\sqrt{p})$ and Lemma 2.7.7 and Corollary 2.8.3. of Vershynin (2019) which provides $\|n^{-1}\sum_{t=1}^{n}u_{jt}f_{t}^{\prime}\|_{\infty}=O_{p}(\sqrt{logp/n})$. Now third term on the right side of (A.104)
@@
$r_{w1}:=\max\left(\sqrt{\frac{logn+\bar{r}}{n\bar{\eta}}},\sqrt{\frac{K}{\bar{\xi}}},\sqrt{\frac{K\bar{\Psi}}{\bar{\xi}\bar{\eta}}},\sqrt{\frac{K}{\bar{\xi}\bar{\eta}}}\right),$ (A.110)
The proof of Theorem 1-Step 2d gives
@@
By (A.113), the third term on the right side of (A.95) is via Theorem 1
$\max_{1\leq j\leq p}\|E(y_{jt}-\mu_{j})(Y_{-jt}-\mu_{-j})^{\prime}\|_{2}\max_{1\leq j\leq p}\|\hat{\alpha}_{j}-\alpha_{j}^{*}\|_{2}=O_{p}(d_{2n}r_{w1}).$ (A.114)
@@
$\max_{1\leq j\leq p}E[(y_{jt}-\mu_{j})(Y_{-jt}-\mu_{-j})^{\prime}]\alpha_{j}^{*}$ $\leq$ $\left[\max_{1\leq j\leq p}\|E[(y_{jt}-\mu_{j})(Y_{-jt}-\mu_{-j})^{\prime}]\|_{2}\right]$ (A.115)
$\times$ $\left[\max_{1\leq j\leq p}\|\alpha_{j}^{*}\|_{2}\right]$